How to find an equation of the line passing through the pair of points in the form of Ax+By=C?
Equation
(-7,8), (-8,-5)


Answer by stanbon(75887)

How to find an equation of the line passing through the pair of points in the form of Ax+By=C?
Equation
(-7,8), (-8,-5)
------------------
slope = (8--5)/(-7--8) = 13/1 = 13
----------------------
Form: y = mx + b
You have m=13 and y = 8 when x = -7 ; solve for "b":
8 = 13*-7 + b
b = 99
---------------
Equation:
y = 13x + 99
-----------------
Rearrange:
13x - y = 99
=================================
Cheers,
Stan H.